# Melding the Serverless Control Plane with the Conventional Cluster Manager for Speed and Resource Efficiency

### v3

@@ 3.2. Analysis of the Control-Plane Scaling Delays
To understand the implications of the workload characteristics we discuss above, we then analyze the delays that occur in the control plane of real-world synchronous and asynchronous Knative-based systems under load. In these experiments, we use the In-Vitro (Ustiugov et al., 2023) load generator that replays a sampled production trace containing 400 functions.
@@
We identify and study two sources of delays in serverless control planes: the instance creation delays and the decision delays. We plot cumulative distribution functions (CDFs) for each source in Figure 3 for vanilla Knative (Kn) and its modified synchronous analogue (Kn-Sync). The figure shows that the sources can cause queuing delays in the range from milliseconds to 10s of seconds, i.e., comparable to the invocation execution time range (Shahrad et al., 2020). Hence, if any of those delays happen on the critical path of invocation handling, they would be seen as noticeable delays to end-to-end request latency. We discuss each delay source in detail below.
3.2.1. Instance Creation Delay.
We find that instance creation takes 1–3s to complete, a delay dominated by features in conventional cluster managers like Kubernetes (Figure 4). The largest contributors include:
Readiness probes introduce a 500ms average delay, a result of the 1-second minimum polling interval Kubernetes uses to protect the control plane at scale.
Namespace and networking setup, requiring multiple round-trips to the cluster manager, adds over 400ms.
Reverse (Queue) Proxy and function sandbox creation contribute over 250ms.
The Golang runtime in the function images we deploy has minimal initialization overhead, but other runtimes like Java can compound this delay by seconds (Ustiugov et al., 2021a).
3.2.2. Decision-Making Delay.
@@
In summary, instance creation delays are the biggest driver of median latency, prompting a redesign of the critical path. While decision-making has a low median latency, its tail must be optimized. An ideal system could achieve ¡200ms end-to-end scaling with fast instance creation, quick decision-making, and no internal congestion.
3.3. Conventional Control Plane Throughput
We evaluate if a conventional control plane can handle the high instance creation rates of FaaS deployments (Shahrad et al., 2020). Using a microbenchmark on a tuned Knative-Kubernetes control plane with emulated worker nodes via KWOK (kwo, 2025), we test its throughput limits (§5).
Our tuned control plane sustains 50 cold starts/sec, a 25 increase over the throughput reported by prior work (Cvetkovic et al., 2024).222This is not a contradiction: our careful configuration and tuning, discussed in §5, has substantially increased Knative control plane’s throughput. (Figure 5). However, when compared to production trace demands, this throughput is 3 lower than the median rate required by an asynchronous system, confirming it can be a bottleneck (Cvetkovic et al., 2024). While sufficient for a synchronous system’s median load, rare 99-th percentile bursts demand 1.2 to 40 more throughput, overwhelming either control plane.
3.4. Cluster Resource Efficiency Analysis
We analyze CPU and memory usage – a major operational cost (Agache et al., 2020) – in our synchronous and asynchronous Knative systems in the same setup as in (§3.2). We find high overheads from two sources: idle instances consume 70% (sync) and 87% (async) of total instance memory, while the control plane’s management tasks, which include activities like instance placement and bootstrapping, consume 9% (sync) and 20% (async) of CPU cycles across the cluster nodes.
This motivates an efficient control plane that provisions instances just-in-time but also minimizes resource waste by reducing instance churn and avoiding a large pool of idle instances with low reuse probability.
3.5. Takeaways
Sustainable and excessive traffic. FaaS traffic is bimodal (§3.1): sustainable traffic consumes the majority (¿98%) of cluster resources with low control plane load, whereas sporadic excessive traffic strains the control plane with instance creation requests despite using few (¡2%) resources. A well-designed system must cater to both.
@@ 4. PulseNet Design
4.1. PulseNet Architecture Overview
@@
Figure 6 shows PulseNet’s workflow when processing invocations and cold starts of new function instances. Function invocations arrive at PulseNet that comprises two components: Load Balancer and Fast Placement. Similar to the existing designs (§2), Load Balancer identifies the target function and finds a Node with an existing non-busy instance of that function to route the invocation for processing. As in the systems with a vanilla asynchronous control plane, Cluster Manager continuously monitors and predicts the trends in the function invocation traffic, adjusting the number of instances for the functions that experience up or down trends in the invocation traffic. We refer to these instances as Regular Instances. Importantly, Regular Instance creations are off the function invocations’ critical path. PulseNet manages Regular Instances similarly to replicas of microservices in conventional managers, such as Kubernetes.
@@
The Load Balancer first tries to route invocations to available Regular Instances. However, if no Regular Instances are available, it marks the remaining invocations as excessive traffic and routes them to the Fast Placement component. Fast Placement triggers an Emergency Instance creation by sending a request to PulseNet’s agent, called Pulselet, on one of the cluster nodes in a Round-Robin. Pulselet spawns an Emergency Instance to process the excessive traffic. Similar to Kubernetes’ kubelet (k8s, 2025a) and Borg’s borglet (Verma et al., 2015), Pulselet manages the lifecycle of the hosting node’s Emergency Instances, shutting each of these instances down when it finishes processing the request it was created for.
4.2. The Expedited Track: Fast, Disposable Emergency Instances
@@
PulseNet’s expedited track achieves speed by spawning Emergency Instances transparently to the cluster manager, limiting the supported feature set, and using checkpoint-restore. PulseNet also lowers the cost by disposing these instances after processing a single invocation.
@@
No cluster manager interaction. Pulselet does not register the Emergency Instance with the conventional cluster manager. This eliminates all associated overheads, including state persistence in etcd, scheduling, and declarative state reconciliation (Cvetkovic et al., 2024). Fast placement periodically retrieves and caches the following information from the cluster manager: function image ID, CPU and memory quotas, and revision information. Since this metadata changes only upon function redeployment or new revision roll-out. Effectively, PulseNet places Emergency Instances in the CPU and memory utilization margin of the cluster, which is always available in real systems as providers tend to keep their server utilization in a 40-80% range (Verma et al., 2015; Tirmazi et al., 2020).
@@
Limited feature set. Emergency Instances support only the features required for processing a single function invocation: OCI container image deployment, outbound network connections via NAT, and locally-enforced CPU and memory quotas (k8s, 2025c), foregoing the features required for long-running, reusable Regular Instances and other microservices. First, for Emergency Instances, PulseNet uses local IP addresses instead of complex CNI integration for service meshes, which requires sending the outbound traffic via the centralized frontend when accessing other services, e.g., storage. Given that Emergency Instances account for a tiny fraction of cluster resources (§3.1), this traffic cannot substantially overload the frontend. Second, Pulselet checks Emergency Instances’ readiness on its own, reducing the load on the centralized cluster manager.
@@
Disposable Instances to Minimize Footprint To prevent resource waste, Emergency Instances are single-use and disposed of after they finish processing an invocation. This ensures that the resources used to handle transient traffic bursts are reclaimed instantly, avoiding the prevalence of idle instances in the existing systems (§3.4).
4.3. The Standard Track: Efficiently Managing Sustainable Traffic with Intelligent Filtering
@@
In practice, we find that the function’s inter-arrival time (IAT) is a good indicator of the trend consistency when comparing it to the keepalive period that the system uses for Regular Instances. Thus, Load Balancer collects and periodically updates the inter-arrival time (IAT) distribution for each function. When the Load Balancer steers an invocation to the expedited track, it also compares the function’s median IAT against PulseNet’s keepalive period. If the median IAT is lower than the keepalive period, it is worth creating an additional long-living Regular Instance for future invocations of that function.333Here, we see the potential for more intelligent mechanisms of metrics filtering, e.g., using a DNN model, but leave it for future work. Hence, Load Balancer includes this invocation into the metrics reported to the cluster manager. Otherwise, the Load Balancer considers the invocation sporadic and filters out the information about it from the metrics.444Note that Load Balancer uses all invocation statistics when collecting IAT distributions. PulseNet’s keepalive and filtering percentile are configurable parameters, which we empirically evaluate in §6.1.
@@
This filtering mechanism prevents the system from creating costly Regular Instances that would likely sit idle, thus reducing the overall memory usage compared to the systems that rely on long keepalive periods (aws, 2025c) or aggressive predictive scaling (Joosen et al., 2023; Roy et al., 2022), as we show in §6.3.3.
4.4. Implementation
@@
We implement PulseNet prototype in vHive (Ustiugov et al., 2021b), an open-source framework widely used for serverless systems research in academia and industry. vHive deploys Knative (kna, 2025a) production FaaS framework, used in commercial serverless offerings (kna, 2025c), atop the Kubernetes (k8s, 2025a) cluster manager. We modify several Knative components, namely Activator and Autoscaler, leaving Kubernetes completely unchanged.
4.4.1. PulseNet Load Balancer and Metric Filtering
@@
We modify the Knative Activator component to serve as a Load Balancer that routes requests to the standard track or expedited track based on the availability of Regular Instances. It also implements the metric filtering as described in §4.3.
4.4.2. Standard Track and Regular Instances
@@
For a thorough evaluation, it is paramount to deploy a production-grade control plane for the PulseNet standard track to have representative overheads present in a real system. Hence, we deploy modified Knative’s control plane that follows the AWS Lambda scaling policy. This control plane runs atop vanilla Kubernetes – with representative interactions across the control plane, cluster manager, and instances.555This approach has a limitation: Knative deploys instances in containers, rather than in MicroVMs like in commercial systems (Agache et al., 2020; web, 2025), which might show shorter sandbox startup latencies. Open-source production-grade MicroVM solutions, e.g., Firecracker or gVisor, come without a production-grade control plane, and re-implementing it in-house with a complete set of features supported by a system like AWS Lambda or Kubernetes might undermine representativity of the control plane interactions.
4.4.3. Expedited Path: Fast Placement, Pulselet, and Emergency Instances
@@
To evaluate PulseNet dual-track design, we need to make sure that our expedited track’s delays are representative of state-of-the-art systems. Hence, we utilize AWS Firecracker MicroVMs as sandboxes for Emergency Instances using their snapshot-restore technology, as in AWS Lambda (Agache et al., 2020; aws, 2025b). For Fast Placement, we use a simple Round-Robin placement algorithm inside Knative Activator to uniformly distribute this load across all worker nodes in the cluster. Emergency Instances run invisible to the cluster manager without tipping the load balance, as they account for a tiny fraction of CPU and memory resources, hence effectively running in the cluster’s overprovisioning margin. We implement Pulselet in Golang, deploying it alongside the regular kubelet.
5. Methodology
@@
Hardware setup. We run PulseNet and other serverless systems on an 8-node c220g5 Cloudlab cluster (Clo, 2025). Each node has two Intel Xeon Silver 4114 CPUs @ 2.20GHz, each with 10 physical cores, 192 GB DRAM, and an Intel SSD.
@@
Real Control-Plane Deployment with an Emulated Large-Scale Cluster. For some experiments in §3 and §6, we use KWOK (kwo, 2025) v0.6.1 to simulate large numbers of worker nodes while retaining a real Knative-Kubernetes control plane, enabling high-scale evaluation of our serverless system without the prohibitive cost of provisioning physical clusters. KWOK supports modeling node and instance behavior to stress-test control plane performance and experiment with scenarios, such as high-rate sandbox creation by the control plane (§3.3) or configurable instance creation time (§6.2.3), which is difficult to observe with real hardware outside a production environment. This approach uses real control plane components with all their interactions, as in a large-scale production deployment, while modeling the load on worker nodes. This makes large-scale and configurable experiments (§6.4.2) feasible in a small-scale research setting.
@@
Software setup. We use vHive (Ustiugov et al., 2021b), i.e., Knative v1.13 running on top of Kubernetes v1.29. Regular Instances run atop containerd v1.6.18, while Emergency Instances execute in AWS Firecracker VMs v1.10.1. We limit each Knative instance’s concurrency to 1 (i.e., maximum per-instance queue depth), similarly to AWS Lambda (aws, 2025c). We assume all container images and VM snapshots are cached in memory on each node, similarly to prior work (Cvetkovic et al., 2024). In all Knative-based configurations to ensure measurement stability, we disable SMT and fix the CPU frequency to the base frequency, disable the Knative panic mode, and set the Load Balancer’s (called Activator in Knative) replica count to 1 but ensure that it never becomes a bottleneck.666During the peak utilization in our experiments, we observe that Knative Activator uses less than one CPU core with a 99-th percentile 10ms routing delay.
@@
Workload. We use In-Vitro (Ustiugov et al., 2023) methodology for representative trace sampling and load generation. In all experiments, we use a 400-function trace sample with per-function IAT and duration distributions, chosen to apply the maximum possible load to the cluster without reaching 100% CPU utilization at any point throughout the experiment. Only in §6.4.2, we use a bigger trace with 2000 functions to evaluate the performance of a larger, 50-node cluster. We run experiments for an hour, discarding the first 20 minutes as a warm-up. Similarly to the prior work (Cvetkovic et al., 2024; Ustiugov et al., 2023), we use a synthetic spin-loop function with a programmed duration taken from the trace.
@@
Baselines. We compare PulseNet to five state-of-the-art serverless systems. We use four baselines based on Knative (kna, 2025a), which is a K8s-based serverless system widely used in commercial offerings (kna, 2025c). First is vanilla Knative (Kn) with an asynchronous control plane, which uses its default concurrency-based autoscaling policy with the default 60-second autoscaling window. We disable panic mode to stabilize the autoscaling. We carefully configure the Knative deployment to maximize the baseline control plane’s performance.777We increase concurrency and request rate limits to the Kubernetes API Server in Knative components and Kubernetes Controller Manager, and also increase the CPU and memory quotas for Knative core components. The second one is Knative-Synchronous (Kn-Sync), the Knative version we implement with synchronous instance creation similar to AWS Lambda. Specifically, we have modified Autoscaler to trigger new instance creations when it cannot find an available instance for a new request and retain the instances for a fixed period of inactivity. We use a 10-minute keep-alive period, which is estimated to be the keep-alive duration for AWS Lambda (Ustiugov et al., 2021a). We also use Dirigent (Cvetkovic et al., 2024), a clean-slate serverless cluster manager with a high-performance asynchronous control plane, which is the state-of-the-art academic system. Finally, (Kn-NHITS) and (Kn-LR) that use NHITS (Challu et al., 2023) and lightweight linear regression prediction models (demonstrated as the most accurate predictors for FaaS by the prior work (Joosen et al., 2023)), respectively, replacing the default Knative autoscaling policy. We train the models on the one-hour-long part of the trace that precedes the part used in the evaluation.
@@
Performance and Cost Metrics. We are using several metrics to evaluate the systems. Our main characteristics of the systems are function invocation performance and the costs that the serverless system incurs to provide that level of performance. To evaluate performance, we use the geometric mean of the tail (99-th percentiles) of per-function slowdown. Lower slowdown is better, with the slowdown of 1 meaning that the system’s end-to-end response time is as low as measured in an unloaded system.888We first calculate per-invocation slowdown by dividing the end-to-end response time by expected execution duration. Then we compute the per-function 99-th percentile of slowdown. Finally, we aggregate per-function slowdown with the geometric mean. To estimate the cost-efficiency, we use the total memory footprint of all instances in the cluster, normalized to the total memory footprint of non-idle instances. We call this metric normalized cost. We also evaluate other cost sources, such as instance creation rates and the CPU cycles used by control plane components. We derive these metrics by collecting cluster-wide metrics with Prometheus (pro, 2025) and Kubernetes Metrics Server (k8s, 2025e).
6. Evaluation

### v4

@@ 3.2 Control-Plane Scaling Delays Analysis
To understand the implications of the workload characteristics we discussed above, we analyze the delays that occur in the control plane of real-world synchronous and asynchronous Knative-based systems under load. In these experiments, we use the In-Vitro [ustiugov:enabling] load generator that replays a sampled production trace containing 400 functions.
@@
We identify and study two sources of delays in serverless control planes: the instance creation delays and the decision delays. We plot cumulative distribution functions (CDFs) for each source in Figure 3 for vanilla Knative (Kn) and its modified synchronous analogue (Kn-Sync). The experiment shows that the component of the queuing delays in the range from milliseconds to 10s of seconds, i.e., comparable to the invocation execution time range [shahrad:serverless]. Hence, if any of those delays happen on the critical path of invocation handling, they would be seen as noticeable delays to end-to-end request latency. We discuss each delay source in detail below.
3.2.1 Instance Creation Delay
We find that instance creation takes 1–3s to complete, a delay dominated by features in conventional cluster managers like Kubernetes (Figure 4). Readiness probes introduce a 500ms average delay, a result of the 1-second minimum polling interval Kubernetes uses to protect the control plane at scale. Namespace and networking setup requires multiple round-trips to the cluster manager, adding over 400ms. Reverse (Queue) Proxy and function sandbox creation contribute over 250ms. Finally, the Golang runtime in the function images we deploy has minimal initialization overhead, but other runtimes like Java can compound this delay by seconds [ustiugov:stellar].
3.2.2 Decision-Making Delay
@@
In summary, instance creation delays are the biggest driver of median latency, prompting a redesign of the critical path. While decision-making has a low median latency, system architects should focus on optimizing its tail. An ideal system could achieve <200ms end-to-end scaling with fast instance creation, quick decision-making, and no internal congestion.
3.3 Conventional Control-Plane Throughput
We evaluate if a conventional control plane can handle the high instance creation rates of FaaS deployments [shahrad:serverless]. Using a microbenchmark on a tuned Knative-Kubernetes control plane with emulated worker nodes via KWOK [kwok], we test its throughput limits (§6).
Our tuned control plane sustains 50 cold starts/sec, a 25 increase over the throughput reported by prior work [cvetkovic:dirigent]222This is not a contradiction: our careful configuration and tuning, discussed in §6, have substantially increased Knative control plane’s throughput. (Figure 5). However, when compared to production trace demands, this throughput is 3 lower than the median rate required by an asynchronous system, confirming it can be a bottleneck [cvetkovic:dirigent]. While sufficient for a synchronous system’s median load, rare 99-th percentile bursts demand 1.2 to 40 more throughput, overwhelming either control plane.
3.4 Cluster Resource Efficiency Analysis
We analyze CPU and memory usage – a major operational cost [agache:firecracker] – in our synchronous and asynchronous Knative systems in the same setup as in (§3.2). We find high overheads from two sources: idle instances consume 70% (sync) and 87% (async) of total instance memory, while the control plane’s management tasks, which include activities like instance placement and bootstrapping, consume 9% (sync) and 20% (async) of CPU cycles across the cluster nodes.
This motivates an efficient control plane that provisions instances just-in-time, but also minimizes resource waste by reducing instance churn and avoiding a large pool of idle instances with low reuse probability.
3.5 Summary and Takeaways
Sustainable and excessive traffic. FaaS traffic is bimodal (§3.1): sustainable traffic consumes the majority (>98%) of cluster resources with low control-plane load, whereas sporadic excessive traffic strains the control plane with instance creation requests despite using few (<2%) resources. A well-designed system must cater to both.
@@ 4 PulseNet Design
4.1 PulseNet Architecture Overview
@@
Figure 6 shows PulseNet’s workflow when processing invocations and cold starts of new function instances. Function invocations arrive at PulseNet that comprises two components: Load Balancer and Fast Placement. Similar to the existing designs (§2), Load Balancer identifies the target function and finds a Node with an existing non-busy instance of that function to route the invocation for processing. As in systems with a vanilla asynchronous control plane, Cluster Manager continuously monitors and predicts trends in function invocation traffic, adjusting the number of instances for functions that experience up or down trends in invocation traffic. We refer to these instances as Regular Instances. Importantly, Regular Instance creations are off the function invocations’ critical path. PulseNet manages Regular Instances similarly to how microservice replicas are managed in conventional managers, such as Kubernetes.
@@
The Load Balancer chooses the track for each request based on the availability of the Regular Instances.333We consider an instance available if its per-instance queue is not full. If there are available Regular Instances in the standard track, Load Balancer uses one of them. Otherwise, it sends the request to expedited track’s Fast Placement component. Fast Placement triggers an Emergency Instance creation by sending a request to PulseNet’s agent, called Pulselet, on one of the cluster nodes in a Round-Robin. Pulselet spawns an Emergency Instance to process the excessive traffic. Similar to Kubernetes’ kubelet [k8s] and Borg’s borglet [verma:borg], Pulselet manages the lifecycle of the hosting node’s Emergency Instances, shutting down each instance once it completes processing the request it was created for.
This track choice rule prioritizes optimally allocated Regular Instances, resorting to Emergency Instances in cases during traffic bursts and fast changes in the traffic trends.
4.2 The Expedited Track: Fast, Disposable Emergency Instances
@@
PulseNet’s expedited track achieves speed by spawning Emergency Instances transparently to the cluster manager, limiting the supported feature set, and using checkpoint-restore. PulseNet also lowers the cost by disposing of these instances after processing a single invocation.
@@
No cluster manager interaction. Pulselet does not register the Emergency Instance with the conventional cluster manager. This eliminates all associated overheads, including state persistence in etcd, scheduling, and declarative state reconciliation [cvetkovic:dirigent]. Fast placement periodically retrieves and caches the following information from the cluster manager: function image ID, CPU and memory quotas, and revision information. Since this metadata changes only upon function redeployment or new revision roll-out.
@@
Limited feature set. Emergency Instances support only the features required for processing a single function invocation: OCI container image deployment, outbound network connections via NAT, and locally-enforced CPU and memory quotas [k8s_quotas], foregoing the features required for long-running, reusable Regular Instances and other microservices. First, for Emergency Instances, PulseNet uses local IP addresses rather than complex CNI integration for service meshes, which requires sending outbound traffic via the centralized frontend to access other services, e.g., storage. Given that Emergency Instances account for a tiny fraction of cluster resources (§3.1), this traffic cannot substantially overload the frontend. Second, Pulselet checks the readiness of Emergency Instances independently, reducing the load on the centralized cluster manager. Note, the features PulseNet supports are enough to execute arbitrary user-function code both in Regular and Emergency Instances.
@@
Disposable Instances to Minimize Footprint To prevent resource waste, Emergency Instances are single-use and disposed of after they finish processing an invocation. This ensures that the resources used to handle transient traffic bursts are reclaimed instantly, avoiding the prevalence of idle instances in the existing systems (§3.4). Emergency Instances process their dedicated invocations to completion and have the same scheduling priority as Regular Instances.
4.3 The Standard Track: Efficiently Managing Sustainable Traffic with Smart Filtering
@@
In practice, we find that the function’s inter-arrival time (IAT) is a good indicator of trend consistency when compared with the keep-alive period used by the system for Regular Instances. Thus, Load Balancer collects and periodically updates the IAT distribution for each function. When the Load Balancer steers an invocation to the expedited track, it also compares the function’s median IAT against PulseNet’s keep-alive period. If the median IAT is lower than the keep-alive period, it is worth creating an additional long-living Regular Instance for future invocations of that function. Hence, Load Balancer includes this invocation into the metrics reported to the cluster manager. Otherwise, the Load Balancer treats the invocation as sporadic and filters it out from the metrics. PulseNet’s keep-alive and filtering percentile are configurable parameters we empirically evaluate in §7.1.
@@
This filtering mechanism prevents the system from creating costly Regular Instances that would likely sit idle, thus reducing the overall memory usage compared to the systems that rely on long keep-alive periods [aws-lambda-scaling] or aggressive predictive scaling [joosen:how, roy:icebreaker], as we show in §7.3.3.
4.4 Discussion
Accuracy of Traffic Classification. Despite the simplicity of our IAT-based filtering technique, it is highly efficient: PulseNet steers the bulk of traffic to Regular Instances (§7.3). Misclassifications are benign: overly conservative filtering may increase the use of Emergency Instances, whereas overly permissive filtering may create more Regular Instances that are recycled by the keep-alive mechanism. We leave the design of a more intelligent traffic classifier to future work.
Resource Allocation. Effectively, PulseNet places Emergency Instances in the CPU and memory utilization margin of the cluster that is typically in the 20-60% range in production [verma:borg, tirmazi:borg] – more than enough to accommodate a mere 10% fraction used by these instances in our experiments (§7.3).
PulseNet Limitations. The PulseNet design relies on the observation that unpredictable bursts comprise a small fraction in production deployments, whereas the majority of function invocations can be tracked with a window-based or predictor-based autoscaler, as we show in §3.1, corroborating prior works [joosen:how, shahrad:serverless]. In the hypothetical scenarios where bursts dominate the traffic, PulseNet’s expedited track might be utilized on par with the standard track, requiring a more sophisticated load balancing policy.
5 PulseNet Implementation
@@
We implement PulseNet prototype in vHive [ustiugov:benchmarking], an open-source framework widely used for serverless systems research in academia and industry. vHive deploys Knative [knative] production FaaS framework, used in commercial serverless offerings [knative_offerings], atop the Kubernetes [k8s] cluster manager. We modify several Knative components, namely Activator and Autoscaler, leaving Kubernetes completely unchanged.
@@
PulseNet Load Balancer and Metric Filtering. We modify the Knative Activator component to serve as a Load Balancer that routes requests to the standard track or expedited track based on the availability of Regular Instances. It also implements the metric filtering as described in §4.3.
@@
Standard Track and Regular Instances. For a thorough evaluation, it is paramount to deploy a production-grade control plane for the PulseNet standard track to have representative overheads present in a real system. Hence, we deploy modified Knative’s control plane that follows the AWS Lambda scaling policy [aws-lambda-scaling] atop vanilla Kubernetes – with representative interactions across the control plane, cluster manager, and instances.444This approach has a limitation: Knative deploys instances in containers, rather than in MicroVMs like in commercial systems [agache:firecracker, web:gvisor], which might show shorter sandbox startup latencies. Open-source production-grade MicroVM solutions, e.g., Firecracker or gVisor, come without a production-grade control plane, and re-implementing it in-house with a complete set of features supported by a system like AWS Lambda or Kubernetes might undermine representativity of the control-plane interactions.
@@
Expedited Path: Fast Placement, Pulselet, and Emergency Instances. To evaluate PulseNet dual-track design, we need to make sure that our expedited track’s delays are representative of state-of-the-art systems. Hence, we utilize AWS Firecracker MicroVMs as sandboxes for Emergency Instances using their snapshot-restore technology, as in AWS Lambda [agache:firecracker, aws-snap-serv]. For Fast Placement, we use a Round-Robin placement algorithm inside Knative Activator to uniformly distribute this load across all worker nodes in the cluster. Emergency Instances run invisibly to the cluster manager without tipping the load balance, as they account for a tiny fraction of CPU and memory resources, hence effectively running in the cluster’s overprovisioning margin. We implement Pulselet in Golang and deploy it alongside the regular kubelet.
6 Methodology
@@
Hardware setup. We run PulseNet and other serverless systems on an 8-node c220g5 Cloudlab cluster [CloudLab]. Each node has two Intel Xeon Silver 4114 CPUs @ 2.20GHz, each with 10 physical cores, 192 GB DRAM, and an Intel SSD.
@@
Real Control-Plane Deployment with an Emulated Large-Scale Cluster. For some experiments in §3 and §7, we use KWOK [kwok] v0.6.1 to simulate large numbers of worker nodes while retaining a real Knative-Kubernetes control plane, enabling high-scale evaluation of our serverless system without the prohibitive cost of provisioning physical clusters. KWOK supports modeling node and instance behavior to stress-test control-plane performance and experiment with scenarios, such as high-rate sandbox creation by the control plane (§3.3) or configurable instance creation time (§7.2.3), which is difficult to observe with real hardware outside a production environment. This approach uses real control-plane components with all their interactions, as in a large-scale production deployment, while modeling the load on worker nodes. This makes large-scale and configurable experiments (§12) feasible in a small-scale research setting.
@@
Software setup. We use vHive [ustiugov:benchmarking], i.e., Knative v1.13 running on top of Kubernetes v1.29. Regular Instances run atop containerd v1.6.18, while Emergency Instances execute in AWS Firecracker VMs v1.10.1, with snapshot-restore enabled. We limit each Knative instance’s concurrency to 1 (i.e., maximum number of concurrent requests per instance), similarly to AWS Lambda [aws-lambda-scaling]. To prove the generality of our approach, we evaluate other values for per-instance concurrency separately in §7.5. We assume all container images and VM snapshots are cached in memory on each node, similarly to prior work [cvetkovic:dirigent]. To ensure measurement stability, we disable SMT and fix the CPU frequency to the base frequency, and set the Load Balancer’s (called Activator in Knative) replica count to 1, but ensure that it never becomes a bottleneck.555During the peak utilization periods in our experiments, Knative Activator uses less than one CPU core with a 99-th percentile 10ms routing delay.
@@
Workload. We use the In-Vitro [ustiugov:enabling] methodology for representative trace sampling and load generation. In all experiments, we use a 400-function trace sample with per-function IAT and duration distributions, chosen to apply the maximum possible load to the cluster without reaching 100% CPU utilization at any point throughout the experiment. Only in §12, we use a bigger trace with 2000 functions to evaluate the performance of a larger, 50-node cluster. We run experiments for an hour, discarding the first 20 minutes as a warm-up. Similar to the prior work [cvetkovic:dirigent, ustiugov:enabling], we use a synthetic spin-loop function with a programmed duration taken from the trace. To confirm our findings with the synthetic functions, we then use a set of realistic serverless functions (encryption, authentication, fibonacci, and image-rotate) from vSwarm [schall:lukewarm], which are written in Go, Python, and Node.js, in §7.5. To achieve the load representative of a production setup, we replicate these functions so that each replica corresponds to one of the functions from the same 400-function sample derived from the Azure traces to fit the sample’s execution duration and memory footprint distributions.
@@
Baselines. We compare PulseNet to five state-of-the-art serverless systems. We use four baselines based on Knative [knative], which is a Kubernetes-based serverless system widely used in commercial offerings [knative_offerings]. First is vanilla Knative (Kn) with an asynchronous control plane, which uses its default concurrency-based autoscaling policy with a 60-second autoscaling window. We carefully configure the Knative deployment to maximize the baseline control plane’s performance.666We increase concurrency and request rate limits to the Kubernetes API Server in Knative components and Kubernetes Controller Manager, and also increase the CPU and memory quotas for Knative core components. The second one is Knative-Synchronous (Kn-Sync), the Knative version we implement with synchronous instance creation similar to AWS Lambda. Specifically, we have modified Autoscaler to trigger new instance creations when it cannot find an available instance for a new request and retain the instances for a fixed period of inactivity. We use a 10-minute keep-alive period, which is estimated to be the keep-alive duration for AWS Lambda [ustiugov:stellar]. We also use Dirigent [cvetkovic:dirigent], a clean-slate serverless cluster manager with a high-performance asynchronous control plane, which is the state-of-the-art academic system. Finally, (Kn-NHITS) and (Kn-LR) that use NHITS [challu:nhits] and lightweight linear regression prediction models (demonstrated as the most accurate predictors for FaaS by the prior work [joosen:how]), respectively, replacing the default Knative autoscaling policy. We train the models on the one-hour-long part of the trace that precedes the part used in the evaluation.
@@
Performance and Cost Metrics. We are using several metrics to evaluate the systems. Our main characteristics of the systems are function invocation performance and the costs that the serverless system incurs to provide that level of performance. To evaluate performance, we use the geometric mean of the tail (99-th percentiles) of per-function slowdown. Lower slowdown is better, with the slowdown of 1 meaning that the system’s end-to-end response time is as low as measured in an unloaded system.777We first calculate per-invocation slowdown by dividing the end-to-end response time by expected execution duration. Then we compute the per-function 99-th percentile of slowdown. Finally, we aggregate per-function slowdown with the geometric mean. To estimate the cost-efficiency, we use the total memory footprint of all instances in the cluster, normalized to the total memory footprint of non-idle instances. We call this metric normalized cost. We also evaluate other cost sources, such as instance creation rates and the CPU cycles used by control-plane components. We derive these metrics by collecting cluster-wide metrics with Prometheus [prometheus] and Kubernetes Metrics Server [k8s-metrics].
7 Evaluation